OpenStudy (anonymous):

HELP!! if f(x) = 20+8/x ; g(x) = 8/x what is (f/g) (8)= ?

OpenStudy (mertsj):

\[f(x)=20+\frac{8}{x}\]

OpenStudy (mertsj):

Is that right?

OpenStudy (anonymous):

No f divided by g then multiplied by 8

OpenStudy (mertsj):

My question is: Do I have the f(x) right? Is it the way I wrote it or is it like this:

OpenStudy (mertsj):

\[f(x)=\frac{20+x}{8}\]

OpenStudy (anonymous):

no the first way you wrote it

OpenStudy (mertsj):

Ok. Replace x with 8. What do you get?

OpenStudy (anonymous):

1

OpenStudy (mertsj):

\[20+\frac{8}{8} = 1???????\]

OpenStudy (anonymous):

8/8 is 1 right

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

20+1=21

OpenStudy (mertsj):

And 20 + 1 is ?????

OpenStudy (mertsj):

Now take the function g which is 8/x and replace x with 8. What do you get?

OpenStudy (anonymous):

8/8=1

OpenStudy (mertsj):

Now put your first answer over your second answer and that will be the f/g(8)

OpenStudy (anonymous):

so is it (21/8) (8) ?